OpenStudy (randomecorner):

Use the distributive property to rewrite the expression and solve. 6 94 = 6 + 6 4 =

OpenStudy (texaschic101):

6 * 94 = 6 * 90 + 6 * 4 = 564

OpenStudy (randomecorner):

how is it all times?

OpenStudy (texaschic101):

yes...6 x 94 is always 564

OpenStudy (randomecorner):

wait never mind *face palms*

OpenStudy (randomecorner):

lol thx

OpenStudy (texaschic101):

lol...anytime :)